QUESTION
TONY HAS GOT HIS KITE STUCK AT THE TOP OF A VERY TALL TREE. HE KNOWS THE STRING ON HIS KITE IS 20 M LONG. WHEN HE PULLS THE STRING TIGHT AND HOLDS THE VERY END ON THE GROUND IT TOUCHES 17 M FROM THE BOTTOM OF THE TREE. IF THE GROUND IS FLAT, HOW TALL IS THE TREE?

Answer by mananth(16946)

the tree ,string & and the ground form a right triangle
string = hypotenuse = 20m
distance from tree =17 m
so height^2 = 20^2-17^2
=400-289
=111
height = sqrt(111)
= 10.5 m height of tree
...
m.ananth@hotmail.ca